A florist brought 240 roses at rs 9 per dozen . if he sold all of them at rs 1 each ,what profit did he make ? Also what is his profit per cent ?

Grade:9

1 Answers

Latika Leekha
9 years ago
Hello student,\
C.P of 1 rose = 9/12 = 0.75
S.P of 1 rose = Re 1
Hence, since S.P. is greater than C.P. so he has made profit.
Profit = S.P. – C.P. = 1-0.75 = 0.25
Hence, profit on sale of 1 rose = 0.25
This means that the profit on sale of 240 roses = 240 x 0.25 = Rs 60
Profit % = (Profit/C.P.) x 100
= (0.25/0.75) x 100
= 100/3